OpenStudy (anonymous):

The connecticut river flows at a ate of 4km/h for the length of a popular scenic route. In order for a cruiser to travel 60km upriver and then return in a total of 8 hr, how fast must the boat be able to travel in still water?

OpenStudy (anonymous):

60 = (v-4)t1 60=(v+4)t2 t1+t2 = 8

OpenStudy (anonymous):

t1 is the time taken upstream, t2 downstream, and v is the speed of the boat

OpenStudy (anonymous):

ok. now what do i do?

OpenStudy (dumbcow):

use series of substitutions to find v t1 = 8-t2 replace 8-t2 for t1 in first equation 60 = (v-4)(8-t2) get t2 by itself in terms of v replace that for t2 in 3rd equation

OpenStudy (anonymous):

uve got three variables nd three eqns substitute t2 and (8-t1) in the first two eqns

OpenStudy (anonymous):

ok.... this is tricky.. lol